Aug 20, 2009 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET and EVO session
  • Phone:
    • +1-800-377-8846 : US
    • +1-888-276-7715 : Canada
    • +1-302-709-8424 : International
    • then enter participant code: 77438230# (remember the "#"). Note: code changed 7/2/09

Items for followup from previous meeting(s)

  1. Review calibration procedures

Documents to Review

Tentative Agenda

  1. Announcements
    1. Report from Bcal Readout Review submitted
  2. Review of Action Items
  3. FCAL Update
    1. Procurement update (Elton)
  4. BCAL Updates
    1. Fiber QA
    2. Lead QA
    3. First-article report draft (Zisis)
    4. Fabrication (Zisis)
      • Reporting
      • Readiness Review ?
  5. Simulations
  6. BCAL Readout
    1. MOU update?
    2. Testing of SiPM arrays (Fernando and Carl)
  7. Other business


Attendees: UofR: Zisis, George, Andrei, Dan; JLab: Tim, John, Beni, Elton; USM: Will, Alam, Manuel, Pedro

  1. Announcements
  2. Review of Action Items
    1. Report from Bcal Readout Review submitted
  3. FCAL Update
    1. Matt is away, Elton reported on procurement status. PR for fabrication of FCAL modules has been placed and in process. Specification for FEU-84-3 pmts has been written and collecting signatures.
  4. BCAL Updates
    1. BCAL Fabrication
      1. 'Big contract' signed (Aug.6). $195,000 until Sept 30, later extended to $220,407. Example.jpg Reporting form received from Teresa.
      2. Ross 'Big Machine' is operational. Big and Small were used and first Al plate delivered a week ago. Slight concave in middle; should be no problem.
      3. Press: 'lunar lander' feet on rams are installed. One outermost cylinder has a leak; parts ordered; does not affect build. New, longer press swivel and locking arms will be made and press retrofitted; end pairs will be made and installed. Media:LunarLander.jpg
      4. Fibre sorting table ready and operational; Cu sheet glued with contact cement and tapped at the ends; grounded.
      5. Vaseline-free construction room. (Steak dinner courtesy of Tim).
      6. Building of the Construction Prototype commenced (Aug. 18). Two builds, almost 20 layers. TasmanianDevil.jpg Tower measurements will be done this morning.
      7. Quotations for 2nd press are under way; contract will most likely be split between Ross and Hydro Designs. Pistons have been tendered.
      8. Danny: Slight crown (6 mils) on current stack. Will continue to monitor.
    2. Fiber QA
      • QA tests on shipment-3 fibres continues. Some irregularities in atten length data: dependence on Kuraray pouch? Data being checked.
      • First-article report draft (Zisis)
      • First-article report checks (Andrei)
      • Will: asked about two components to attenuation. Zisis said fits have been made and will be made available.
    3. Lead QA
    4. Reporting excel sheet has been developed and must be filled out by UofR once per month in advance of project reporting.
    5. Readiness Review. Tentative schedule for first week of October.
  5. Simulations
    • No report
  6. BCAL Readout
    • Light guide simulations continue and construction of dark box. Report at next week.
  7. Other business